### v1

@@ 3.1 Proposed Method
The study of maximum misclassification rates in Section 2.2 answers the question of when misclassification bias is negligible in the two-stage framework. However, it is still unclear how the desired misclassification rate in Section 2.2 can be achieved in the two-stage framework. Although various two-stage approaches have been proposed (see an overview in bakk2021relating), they often lack theoretical guarantees, especially in high dimensions, and might not be able to achieve the desired misclassification rate to infer latent subgroup effects consistently and efficiently (bakk2014relating). Specifically, existing two-stage approaches typically fit the latent class model via the Expectation-Maximization (EM; dempster1977maximum) algorithm. Such likelihood-based procedures lack algorithmic convergence guarantees and can suffer from initialization difficulties and large computational cost, especially in the presence of a large number of items (kwon2019global; balakrishnan2017statistical).
Indeed, zeng2023tensor showed that instead of being viewed as a challenge, the increasing availability of a relatively large number of item responses in the latent class model can help decrease the misclassification rate if the latent class model is estimated appropriately. To utilize the large number of item responses to achieve the desired misclassification rate, we introduce the spectral method perspective into the two-stage framework and particularly, adapt the spectral clustering to estimate the latent subgroup memberships in the first stage. Different from the likelihood-based method, spectral clustering groups the subjects by leveraging the approximate low-rank structure of the response data matrix and thus has the potential to efficiently accommodate high-dimensional data and capture their valuable information. See a comprehensive monograph on spectral methods in chen2021spectral and references therein. Spectral clustering has been extensively studied and adopted in various scenarios, such as sub-Gaussian mixture models and network models (abbe2022; srivastava2023robust; zhang2022leave; li2023community), but it has not been investigated in the latent subgroup setting considered here. Although recent studies in lyu2024degree and lyu2025spectral examined spectral clustering in latent class models, they did not consider the statistically challenging downstream inference on the latent subgroup effects. To estimate and infer the latent subgroup effect $\mu$ in a valid and efficient way, we next propose a spectral-based two-stage approach as summarized in Algorithm 1. For simplicity of presentation, we rewrite the subgroup outcome model Eq. (1) in matrix notation with $\mathbf{Y}=\{Y_{i}\}_{i=1}^{n}$, $\mathbf{D}=\text{Diag}\{D_{i}\}_{i=1}^{n}$ and $\mathbf{X}=(X_{1}^{\mathsf{T}},\ldots,X_{n}^{\mathsf{T}})^{\mathsf{T}}$.
@@
In Algorithm 1, we first project the item responses $\mathbf{R}$ into a low-dimensional space of dimension $G$ via top-G SVD decomposition in step 1, and then estimate the latent group membership $\widehat{Z}$ via K-means++ based on the projected item response in step 2. In the end, we proceed to the second stage estimation or inference by regressing $Y$ on $\widehat{Z},D\widehat{Z},X$. The number of latent subgroups $G$ can be determined by prior knowledge or parallel analysis (dobriban2019deterministic). By leveraging the approximate low-rankness of item response matrix $\mathbf{R}$, we reduce the dimension from $J$ to $G$ before performing the K-means++ clustering. Since typically $G\ll J$, such a spectral-based approach substantially reduces the computational burden and alleviates the issue of model misspecification and local optima while preserving the useful latent information. This helps achieve the desired misclassification rate with the blessing of many item responses as detailed in Section 3.2.
@@ 3.2 Theoretical Justification
To justify the proposed two-stage spectral subgroup analysis in Algorithm 1, we start with an introduction of ($J_{I}$$\delta$)-separability adapted from najafi2020reliable, a key concept measuring the separation across different subgroup centers. In specific, by Definition 1, ($J_{I}$$\delta$)-separability requires that among all the $J$ items in the study, there are at least $J_{I}$ informative items whose item parameters are well separated by at least the amount of $\delta$ across different subgroups. Clearly, a larger $J_{I}$ or $\delta$ indicates more information contained in the item responses.
@@
To justify our method, besides the signal-to-noise ratio, Assumptions 2 and 3 introduce two mild requirements on balanced subgroups and singular values, respectively. Assumption 2 states that the sizes of latent subgroups should be balanced and is easily satisfied in practice, such as congenital depressed and acquired depressed patient subgroups studied in loffler2021optimality. Assumption 3 states that the singular values of the item parameter matrix have a similar magnitude, which is a reasonable assumption under common data generative schemes in latent class models following lyu2024degree. For example, Assumption 3 is satisfied with high probability when $\theta_{j,g}\overset{i.i.d.}{\sim}U[0,1]$ for all $j\in[J]$ and $g\in[G]$. Under Assumptions 1-3, we show how the misclassification rate decays with respect to the number of informative items $J_{I}$ in Theorem 3.
@@
From Table 1, we note that as expected, the design conditions are more stringent in high dimensions than low dimensions, and more stringent to achieve valid inference compared to consistent estimation. All design conditions are easily satisfied in practice, as even the most stringent one, that for high-dimensional inference, only requires the number of informative items $J_{I}$ to increase at a rate faster than $\log(n)$ when the separability $\delta$ is a constant. Moreover, because the desired misclassification rates in Theorems 1 and 2 are sharp and the exponential rate in Theorem 3 is optimal for the mixture of isotropic Gaussians under signal-to-noise conditions as discussed in zhang2022leave, to this extent, the desired study design conditions in Table 1 are nearly optimal.
@@ 4 Simulation Studies
In this section, we conduct simulation studies to assess the finite sample performance of the proposed two-stage spectral latent subgroup analysis. We start with a scenario where all items are informative, and consider $G=3$ latent subgroups with the total sample size $n=150$. In the latent class model, we use a similar setup as lyu2025spectral and consider balanced latent class proportions; i.e., $\mathbb{P}(Z=g)=1/G$. We generate the (informative) item responses $\mathbf{R}$ by Eq. (2) with the well separated item parameter matrix $\boldsymbol{\Theta}$, which vertically stacks $J/5$ copies of the following $5\times 3$ base matrix $\boldsymbol{\Theta}_{b}$:
@@
We vary the number of (informative) items across $J_{I}=J=5,10,20,\ldots,150$. In the subgroup outcome model, we follow the data generation in shinkyu2025small, and consider the $p$-dimensional potential confounders $X_{i}\sim N(0,\mathbf{\Sigma})$ with $p=10$ and $\mathbf{\Sigma}_{j,k}=0.1^{|j-k|}$ for the low-dimensional setting and $p=200$ for the high-dimensional setting. The treatment indicator $D$ follows logistic distribution $\mathbb{P}(D=1|X)=1/(1+\exp\{-(1+X^{\mathsf{T}}\beta_{l})\})$ with coefficient $\beta_{l}=(1,1,1,1,1,1,0,\ldots,0)$, and the outcome of interest $Y$ is generated by Eq. (1) with the latent subgroup treatment effect $\mu=(-1,0,1.5)$, main effect $\alpha=(0,-0.05,-0.2)$, sparse regression coefficient $\beta=(1,1.5,1.5,0,\ldots,0)^{\mathsf{T}}$ and random error $\varepsilon\sim N(0,1)$.
We implement the proposed two-stage spectral latent subgroup analysis with the algorithm as detailed in Section 3. In particular, we assume $G$ is known, and adopt ordinary least squares for both estimation and inference in low dimensions. We adopt Lasso and debiased Lasso with cross-validated tuning parameters for estimation and inference, respectively, in high dimensions (shinkyu2025small). For comparison, we also implement the following methods:
@@ 5 Application to TIMSS Educational Assessment Data
In this section, we apply our method to analyze a dataset from the Trends in International Mathematics and Science Study (TIMSS). TIMSS is an observational study investigating the performance of students in mathematics and science at the fourth and eighth grades in more than 50 participating countries and areas around the world since 1995. Besides students’ achievement scores, the TIMSS has collected via questionnaires a rich array of students’ demographic information and students’ agreement levels to statements of motivation for learning mathematics and science, and educational programs students attended. Over the past decades, TIMSS has been widely used to monitor curricular implementation and identify promising instructional practices. In particular, lyu2023estimating identified that private science lessons might have heterogeneous effects on achievement scores of science in the latent subgroups of students with different motivations. However, such analyses using Bayesian joint estimation are based on the one-stage framework of latent subgroup analysis with observational data, which might be inadequate to adjust for the high-dimensional potential confounders.
@@
Here, we focus on 2271 eighth-grade students of TIMSS from Hong Kong in 2015 and aim to uncover among these students, how private science lessons impact the science assessment performance of students with different motivations in science and mathematics. We let the treatment indicator $D$ denote whether or not a student has attended private science lessons during the last 12 months. Let the outcome of interest $Y$ denote the scaled scores of students’ science achievement. The latent subgroups of interest $Z$ are student subgroups with different motivations. The item responses $R$ are the binary answers (agree or disagree) to $J=73$ statements about motivations for learning mathematics and science in the collected questionnaires. The potential confounders $X$ consist of the demographic information and the binary answers to statements, and their interaction terms. For the answers to statements in potential confounders, we follow the practice in lyu2023estimating and assign scores of 0, $\frac{1}{3}$, $\frac{2}{3}$ and $1$ to the answers ”Disagree a lot”, ”Disagree a little”, ”Agree a little” and ”Agree a lot”, respectively to quantify the students’ attitude to science and mathematics. The total number of potential confounders in this study is high-dimensional with $p=6676$. Previous studies of the TIMSS are typically based on the one-stage framework and thus can not accommodate many confounders. For example, lyu2023estimating only includes 18 covariates in their study. As discussed in pearl2022comment, a limited number of confounders might hinder the interpretation of study conclusions. Here, in the presence of high-dimensional confounders, we implement the proposed two-stage spectral latent subgroup analysis in the same way as that for high dimensions in Section 4 and choose the number of latent subgroups by parallel analysis (dobriban2019deterministic).
@@ 6 Discussion
In this paper, we consider a widely seen two-stage scenario where latent subgroup memberships are treated as covariates in the subgroup outcome model. In practice, latent subgroup memberships might be treated as the outcome of interest in the subgroup outcome model, particularly in many classification problems. For example, political scientists are interested in using demographic covariates to predict subgroups of ideological orientations inferred from survey responses (wright2023beyond), and market analysts are interested in using business records to predict subtypes of consumers inferred from purchase patterns (wedel2000market). When latent subgroup membership is treated as the outcome of interest, it is important to understand when and how the two-stage framework remains valid. We leave this interesting direction for future studies.
7 Acknowledgments
The authors thank the anonymous reviewers for their valuable suggestions. YL and XG is partially supported by a grant from Research Grants Council of the Hong Kong Special Administrative Region, China (HKUST 26308323), the Seed fund of the Big Data for Bio-Intelligence Laboratory (Z0428) and the grant L0438 from the Hong Kong University of Science and Technology. YG is partially supported by NSF grant DMS-2210796.

### v2

@@ 3.1 Proposed Method
The study of maximum misclassification rates in Section 2.2 answers the question of when misclassification bias is negligible in the two-stage framework. However, it is still unclear how the desired misclassification rate in Section 2.2 can be achieved in the two-stage framework. Although various two-stage approaches have been proposed (see an overview in Bakk and Kuha,, 2021), they often lack theoretical guarantees, especially in high dimensions, and might not be able to achieve the desired misclassification rate to infer latent subgroup effects consistently and efficiently (Bakk et al.,, 2014). Specifically, existing two-stage approaches typically fit the latent class model via the Expectation-Maximization (EM; Dempster et al., (1977)) algorithm. Such likelihood-based procedures lack algorithmic convergence guarantees and can suffer from initialization difficulties and large computational cost, especially in the presence of a large number of items (Kwon et al.,, 2019; Balakrishnan et al.,, 2017).
Indeed, Zeng et al., (2023) showed that instead of being viewed as a challenge, the increasing availability of a relatively large number of item responses in the latent class model can help decrease the misclassification rate if the latent class model is estimated appropriately. To utilize the large number of item responses to achieve the desired misclassification rate, we introduce the spectral method perspective into the two-stage framework and particularly, adapt the spectral clustering to estimate the latent subgroup memberships in the first stage. Different from the likelihood-based method, spectral clustering groups the subjects by leveraging the approximate low-rank structure of the response data matrix and thus has the potential to efficiently accommodate high-dimensional data and capture their valuable information. See a comprehensive monograph on spectral methods in Chen et al., (2021) and references therein. Spectral clustering has been extensively studied and adopted in various scenarios, such as sub-Gaussian mixture models and network models (Abbe et al.,, 2022; Srivastava et al.,, 2023; Zhang and Zhou,, 2024; Li et al.,, 2023), but it has not been investigated in the latent subgroup setting considered here. Although recent studies in Lyu et al., (2025) and Lyu and Gu, (2025) examined spectral clustering in latent class models, they did not consider the statistically challenging downstream inference on the latent subgroup effects. To estimate and infer the latent subgroup effect $\mu$ in a valid and efficient way, we next propose a spectral-based two-stage approach as summarized in Algorithm 1. For simplicity of presentation, we rewrite the subgroup outcome model Eq. (1) in matrix notation with $\mathbf{Y}=\{Y_{i}\}_{i=1}^{n}$, $\mathbf{D}=\text{Diag}\{D_{i}\}_{i=1}^{n}$ and $\mathbf{X}=(X_{1}^{\mathsf{T}},\ldots,X_{n}^{\mathsf{T}})^{\mathsf{T}}$.
@@
In Algorithm 1, we first project the item responses $\mathbf{R}$ into a low-dimensional space of dimension $G$ via top-G SVD decomposition in step 1, and then estimate the latent group membership $\widehat{Z}$ via K-means++ based on the projected item response in step 2. In the end, we proceed to the second stage estimation or inference by regressing $Y$ on $\widehat{Z},D\widehat{Z},X$. The number of latent subgroups $G$ can be determined by prior knowledge or parallel analysis (Dobriban and Owen,, 2019). By leveraging the approximate low-rankness of item response matrix $\mathbf{R}$, we reduce the dimension from $J$ to $G$ before performing the K-means++ clustering. Since typically $G\ll J$, such a spectral-based approach substantially reduces the computational burden and alleviates the issue of model misspecification and local optima while preserving the useful latent information. This helps achieve the desired misclassification rate with the blessing of many item responses as detailed in Section 3.2.
@@ 3.2 Theoretical Justification
To justify the proposed two-stage spectral subgroup analysis in Algorithm 1, we start with an introduction of ($J_{I}$$\delta$)-separability adapted from Najafi et al., (2020), a key concept measuring the separation across different subgroup centers. In specific, by Definition 1, ($J_{I}$$\delta$)-separability requires that among all the $J$ items in the study, there are at least $J_{I}$ informative items whose item parameters are well separated by at least the amount of $\delta$ across different subgroups. Clearly, a larger $J_{I}$ or $\delta$ indicates more information contained in the item responses.
@@
To justify our method, besides the signal-to-noise ratio, Assumptions 2 and 3 introduce two mild requirements on balanced subgroups and singular values, respectively. Assumption 2 states that the sizes of latent subgroups should be balanced and is easily satisfied in practice, such as congenital depressed and acquired depressed patient subgroups studied in Löffler et al., (2021). Assumption 3 states that the singular values of the item parameter matrix have a similar magnitude, which is a reasonable assumption under common data generative schemes in latent class models following Lyu et al., (2025). For example, Assumption 3 is satisfied with high probability when $\theta_{j,g}\overset{i.i.d.}{\sim}U[0,1]$ for all $j\in[J]$ and $g\in[G]$. Under Assumptions 1-3, we show how the misclassification rate decays with respect to the number of informative items $J_{I}$ in Theorem 3.
@@
From Table 1, we note that as expected, the design conditions are more stringent in high dimensions than low dimensions, and more stringent to achieve valid inference compared to consistent estimation. All design conditions are easily satisfied in practice, as even the most stringent one, that for high-dimensional inference, only requires the number of informative items $J_{I}$ to increase at a rate faster than $\log(n)$ when the separability $\delta$ is a constant. Moreover, because the desired misclassification rates in Theorems 1 and 2 are sharp and the exponential rate in Theorem 3 is optimal for the mixture of isotropic Gaussians under signal-to-noise conditions as discussed in Zhang and Zhou, (2024), to this extent, the desired study design conditions in Table 1 are nearly optimal.
@@ 4 Simulation Studies
In this section, we conduct simulation studies to assess the finite sample performance of the proposed two-stage spectral latent subgroup analysis. We start with a scenario where all items are informative, and consider $G=3$ latent subgroups with the total sample size $n=150$. In the latent class model, we use a similar setup as Lyu and Gu, (2025) and consider balanced latent class proportions; i.e., $\mathbb{P}(Z=g)=1/G$. We generate the (informative) item responses $\mathbf{R}$ by Eq. (2) with the well separated item parameter matrix $\boldsymbol{\Theta}$, which vertically stacks $J/5$ copies of the following $5\times 3$ base matrix $\boldsymbol{\Theta}_{b}$:
@@
We vary the number of (informative) items across $J_{I}=J=5,10,20,\ldots,150$. In the subgroup outcome model, we follow the data generation in Shinkyu and Sueishi, (2025), and consider the $p$-dimensional potential confounders $X_{i}\sim N(0,\mathbf{\Sigma})$ with $p=10$ and $\mathbf{\Sigma}_{j,k}=0.1^{|j-k|}$ for the low-dimensional setting and $p=200$ for the high-dimensional setting. The treatment indicator $D$ follows logistic distribution $\mathbb{P}(D=1|X)=1/(1+\exp\{-(1+X^{\mathsf{T}}\beta_{l})\})$ with coefficient $\beta_{l}=(1,1,1,1,1,1,0,\ldots,0)$, and the outcome of interest $Y$ is generated by Eq. (1) with the latent subgroup treatment effect $\mu=(-1,0,1.5)$, main effect $\alpha=(0,-0.05,-0.2)$, sparse regression coefficient $\beta=(1,1.5,1.5,0,\ldots,0)^{\mathsf{T}}$ and random error $\varepsilon\sim N(0,1)$.
We implement the proposed two-stage spectral latent subgroup analysis with the algorithm as detailed in Section 3. In particular, we assume $G$ is known, and adopt ordinary least squares for both estimation and inference in low dimensions. We adopt Lasso and debiased Lasso with cross-validated tuning parameters for estimation and inference, respectively, in high dimensions (Shinkyu and Sueishi,, 2025). For comparison, we also implement the following methods:
@@ 5 Application to TIMSS Educational Assessment Data
In this section, we apply our method to analyze a dataset from the Trends in International Mathematics and Science Study (TIMSS). TIMSS is an observational study investigating the performance of students in mathematics and science at the fourth and eighth grades in more than 50 participating countries and areas around the world since 1995. Besides students’ achievement scores, the TIMSS has collected via questionnaires a rich array of students’ demographic information and students’ agreement levels to statements of motivation for learning mathematics and science, and educational programs students attended. Over the past decades, TIMSS has been widely used to monitor curricular implementation and identify promising instructional practices. In particular, Lyu et al., (2023) identified that private science lessons might have heterogeneous effects on achievement scores of science in the latent subgroups of students with different motivations. However, such analyses using Bayesian joint estimation are based on the one-stage framework of latent subgroup analysis with observational data, which might be inadequate to adjust for the high-dimensional potential confounders.
@@
Here, we focus on 2271 eighth-grade students of TIMSS from Hong Kong in 2015 and aim to uncover among these students, how private science lessons impact the science assessment performance of students with different motivations in science and mathematics. We let the treatment indicator $D$ denote whether or not a student has attended private science lessons during the last 12 months. Let the outcome of interest $Y$ denote the scaled scores of students’ science achievement. The latent subgroups of interest $Z$ are student subgroups with different motivations. The item responses $R$ are the binary answers (agree or disagree) to $J=73$ statements about motivations for learning mathematics and science in the collected questionnaires. The potential confounders $X$ consist of the demographic information and the binary answers to statements, and their interaction terms. For the answers to statements in potential confounders, we follow the practice in Lyu et al., (2023) and assign scores of 0, $\frac{1}{3}$, $\frac{2}{3}$ and $1$ to the answers ”Disagree a lot”, ”Disagree a little”, ”Agree a little” and ”Agree a lot”, respectively to quantify the students’ attitude to science and mathematics. The total number of potential confounders in this study is high-dimensional with $p=6676$. Previous studies of the TIMSS are typically based on the one-stage framework and thus can not accommodate many confounders. For example, Lyu et al., (2023) only includes 18 covariates in their study. As discussed in Pearl, (2022), a limited number of confounders might hinder the interpretation of study conclusions. Here, in the presence of high-dimensional confounders, we implement the proposed two-stage spectral latent subgroup analysis in the same way as that for high dimensions in Section 4 and choose the number of latent subgroups by parallel analysis (Dobriban and Owen,, 2019).
@@ 6 Discussion
In this paper, we consider a widely seen two-stage scenario where latent subgroup memberships are treated as covariates in the subgroup outcome model. In practice, latent subgroup memberships might be treated as the outcome of interest in the subgroup outcome model, particularly in many classification problems. For example, political scientists are interested in using demographic covariates to predict subgroups of ideological orientations inferred from survey responses (Wright and Volodarsky,, 2023), and market analysts are interested in using business records to predict subtypes of consumers inferred from purchase patterns (Wedel and Kamakura,, 2000). When latent subgroup membership is treated as the outcome of interest, it is important to understand when and how the two-stage framework remains valid. We leave this interesting direction for future studies.